Determine the value of π‘Ž so that 𝑨 = 2π’Š βˆ’ π‘Žπ’‹ + π’Œ and 𝑩 = 4π’Š βˆ’ 2𝒋 βˆ’ 2π’Œ are perpendicular.


1
Expert's answer
2022-06-10T09:03:31-0400

"A=(2,-a,1)", "B=(4,-2,-2)".


Two vectors are perpendicular iff "A\cdot B=0".


"A\cdot B=2\cdot4 + (-a)\cdot(-2) + 1\cdot(-2)=8+4a-2=4a+6".


"4a+6=0", so "a=-\frac32".